1/1

Eugen Roth,  Eugen Roth,
Das Kursbuch  La fervoja horaro

translated by Manfredo Ratislavo

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Author of this German poem is Eugen Roth (∗1895-01-24 – †1976-04-28).

Arg-88-180 (2004-02-01 18:36:35)

Translation of the German poem “Das Kursbuch” by Eugen Roth (∗1895-01-24 – †1976-04-28) into Esperanto by Manfredo Ratislavo (Manfred Retzlaff, Stettiner Str. 16, D-59302 Oelde, Germanio, ∗1938-11-04) on 1980-11.

MR-004-2 / Arg-88-181 (2004-02-01 18:36:22)

http://www.poezio.net/version?poem-id=88&version-id=180,181